Thirty animals are to be used in a medical experiment on diet deficiency: 3 male and 7 female rhesus monkeys, 6 male and 4 female chimpanzees, and 2 male and 8 female dogs. If one animal is selected at random, what is the probability of getting chimpanzee or a male?
Answer #1

Answer)

Probability is given by favorable/total.

Total = 30.

Favorable = chimpanzee or a male = 3 + 6 + 4 + 2 = 15.

Probability = 15/30 = 0.5